Questions 1. Using your experimental data for e/m and equation (4) compute the velocity of the electrons for each value of the accelerating potential V. Discuss the range of the velocities the electrons acquired in this experiment. 2. What would be the theoretical slone ofe oranh of Vas a netion o 2

media%2Fee2%2Fee21b88a-f36d-417a-9292-92

media%2Fce1%2Fce14f8f4-809e-41a1-a9bc-32
Answer #1

v=\sqrt {\frac {2eV}{m}}

\frac {e}{m}=1.76\times 10^{11}\: C/kg

v= 5.93\times 10^5 \sqrt V

Put the values of potential V from your experimental data to get the values of velocities for different potentials.